OpenStudy (anonymous):

Simplify -6+[4-(x-2)]

OpenStudy (lalaly):

-6+4-x+2 = -x

OpenStudy (anonymous):

Hello (: make sure to use PEMDAS. parentheses, exponents, multiplication, then addition and finally subtract.

OpenStudy (anonymous):

your answer would be -6=4-x-2+=-x (: